OpenStudy (anonymous):

Total power generated by wind worldwide doubles every 3 years. In a particular year, the world wind-generating capacity was about 90 thousand megawatts. Find the continuous growth rate as a percent and give a formula for wind generating capacity W (in thousand megawatts) as a function of t, the number of years in the future.

OpenStudy (anonymous):

I'm not sure what I am doing wrong... So far, I have... W(t)=ae^kt W(3)=90e^3k 180=90e^3k 2=e^3k ln(2)/3 = k .231 = k So, 23% ? & that would make my equation W(t)=90e^(.231t)

OpenStudy (dan815):

ok

OpenStudy (dan815):

at what growth rate would it be doubling every 3 years

OpenStudy (dan815):

growth rate per year

OpenStudy (dan815):

x(1+r)^3=2x so r=2^(1/3) -1

OpenStudy (dan815):

now we know the growth rate per year we are dealing with

OpenStudy (dan815):

dy/dx= r*y

OpenStudy (dan815):

alll that is left is to solve this differential equation

OpenStudy (dan815):

and solve it for our initial condition

OpenStudy (dan815):

@ganeshie8 does this work?

OpenStudy (dan815):

dont worry Ganeishie will check it for us, let me try to sovle completely meanwhile

OpenStudy (dan815):

oh i cant use that 2^(1/3)-1 formula, because this s contrinous!

OpenStudy (dan815):

y=90e^(.231t), is right

OpenStudy (dan815):

dy/dx=ky dy/y=kdx ln y= kx + c y=ce^kx y(0)=90 setting this gives c=90 y=90*e^kx y(3)=180=90e*3k k=ln(2)/3